# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v4

@@ Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are serial computing, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning models go deeper. Therefore, this paper proposes a deep learning parallelization strategy based on the Universal Approximation Theorem (UAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
@@ 1 Introduction
Currently, deep learning is advancing towards larger models and datasets, driven by the success of ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), Llama (Touvron et al. 2023), and ViT (Dosovitskiy et al. 2020), among others. These large models demonstrate significantly enhanced performance compared to their predecessors. According to UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024b), residual-based convolutional neural networks (CNNs) and Transformer-based large language models are specific implementations of the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989), capable of dynamically approximating functions based on inputs. From the perspective of UAT, larger models inherently possess stronger dynamic fitting capabilities (Wang and Li 2024b, a). Therefore, the trend towards larger models is undoubtedly the future direction of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexities.
@@
Current solutions include model optimization techniques such as quantization (Jacob et al. 2017; Lu, Zhong, and Yang 2022; Guo, Wang, and Cui 2021; Lin et al. 2023), pruning (Ma, Fang, and Wang 2023; Xia et al. 2023), and knowledge distillation (Gou et al. 2020; Gu et al. 2021). Additionally, specialized accelerators (Jouppi et al. 2017; Elbtity, Chandarana, and Zand 2024) and edge computing (Shi et al. 2016; Gale et al. 2020), as well as parallel and distributed computing (Shallue et al. 2018; Ben-Nun and Hoefler 2018; Sergeev and Balso 2018), offer potential solutions. However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
@@
The root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data. Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet (Krizhevsky, Sutskever, and Hinton 2012) and VGG (Simonyan and Zisserman 2014). The introduction of ResNe t (He et al. 2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input (Wang and Li 2024b, a), influencing subsequent network designs and establishing the tradition of multi-layer stacking. While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
@@
Therefore, it is imperative to explore parallelization techniques for deep learning networks, enabling parallel computation across multiple layers. This approach is key to accelerating the inference process, breaking the time bottleneck, and advancing computer vision and natural language processing technologies towards greater efficiency and speed. To develop such parallel computation, we must start from the fundamental theory of deep learning: the UAT. The network must be designed in accordance with UAT principles and must satisfy the requirements outlined in the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), which emphasize the ability to dynamically fit functions based on data. With this in mind and leveraging existing network structures, we designed a parallel network, Para-Former, to validate the feasibility of our theory. Our contributions include:
  • Proposing the theory of parallel deep learning networks and experimentally validating its feasibility.
@@
  • Analyzing the characteristics, advantages, and disadvantages of both serial and parallel networks.
@@
  • Investigating the issues present in deep learning from a data perspective.
The structure of this paper is as follows: In Section 2, we first present the general form of the UAT. Based on the mathematical form of UAT, we define the parallel network Para-Former in Section 3. In Section 4, we validate the effectiveness of Para-Former in various ways: we assess its feasibility in Section 4.2, explore the impact of network depth in Section 4.3, and analyze the influence of datasets on network performances in Section 4.4.
2 UAT
Deep learning networks have demonstrated their powerful capabilities in numerous CV and NLP tasks. In the papers UAT2LLMs (Wang and Li 2024b) and UAT2CVs (Wang and Li 2024a), they have been proven that both Transformer-based networks and residual-based CNNs are specific implementations of the UAT. Therefore, we consider UAT to be the fundamental theory for deep learning. To design parallelized networks, we must start from the principles of UAT. Before formally designing parallel networks, we will first review the basic form of the UAT theory and then propose a network parallelization solution based on this mathematical theory. The UAT was initially proposed by Cybenko (Cybenko 2007), and its basic form is as follows:
@@ 2 UAT
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White  (Hornik, Stinchcombe, and White 1989) further proved that this function can approximate any Borel measurable function. Observing Eq. (1), where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Eq. (1): the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
Figure 1: The general description of the serial network and parallel network.
3 UAT for Parallel Network
3.1 Requirements for Basic Block in Parallel Network
@@
From Eq. (1), we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the parameters between different layers do not interact with each other. Therefore, based on Eq. (1), we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula of $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. (1), the network will face a limitation once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. So to ensure the network can dynamically approximate functions based on input, some or all of the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ must be influenced by the input (referencing to UAT2LLMs and UAT2CVs), which can be simply written as $\alpha_{j}(\mathbf{x})$, $W_{j}(\mathbf{x})$ and $\theta_{j}(\mathbf{x})$. This allows the network to dynamically adapt and fit the function based on the input while enabling parallel computation. So we could design the basic block in Figure 1, like $\text{Block}_{1}\cdots\text{Block}_{n}$ are the implementation of $\alpha_{j}(\mathbf{x})\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}(\mathbf{x})\mathbf{x}+\theta_{j}(\mathbf{x})\right),i=1\cdots n$. Some or all the parameters $\alpha_{j}$, $W_{j}$ and $\theta_{j}$ will change according to input.
@@
3.2 Para-former
Based on the previous request on the basic module, we designed the Para-Former which is a parallel network by using the Transformer module as the basic module. We chose Transformer for this design because it is a well-established and widely-used network architecture, excelling in both CV and NLP. Another reason is that its basic mathematical form aligns with the requirements of Section 3.1.
Before formally introducing the structure of Para-Former, let’s briefly review the Transformer structure. The basic structure of a Transformer is shown in Figure 2, and for convenience, we refer to it as a Transformer module. Typically, it is represented as follows:
$\mathring{\mathbf{H}}=$ $\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (3)
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{i1},\ldots,\hat{\mathbf{H}}_{ih}\right)\mathbf{W}_{O}$
$\hat{\mathbf{H}}_{ij}$ $=\operatorname{Attention}(\mathbf{x}_{ij}\mathbf{W}_{Qij}=\mathbf{Q}_{ij},\mathbf{x}_{ij}\mathbf{W}_{Kij}=\mathbf{K}_{ij},$ (4)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{ij})$
$=\operatorname{softmax}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{T}}{\sqrt{M}}\right)\mathbf{V}_{ij}$
$=\mathbf{H}_{ij}\mathbf{V}_{ij}=\mathbf{H}_{ij}[\mathbf{x}_{ij}\mathbf{W}_{Vij}]$
where the subscript of $ij$ represents the $j$-th head of Multi-Head Attention (MHA) in the $i$-th layer. Please note that throughout the following text, all the subscript $i$ refer to the layer index in the network, like $\mathbf{x}_{i}$. According to Figure 2, assuming the input to the Feed-Forward Network (FFN) in the $i$-th layer of the Transformer, it can be expressed as:
$FFN(\mathbf{x}_{i})=\mathbf{W}_{i,3}\sigma(\mathbf{W}_{i,2}\mathbf{x}_{i}+\mathbf{b}_{i,2})+\mathbf{b}_{i,3}\$ (5)
The above representation mainly describes the engineering process rather than its true mathematical form. According to UAT2LLMs, the MHA and FFN components of Transformers can be expressed in matrix-vector form. We denote the matrix-vector form of variables and operations with a prime symbol (e.g., $\mathbf{x}_{i}$ becomes $\mathbf{x}^{\prime}_{i}$). So MHA and FFN can be written as (For more details, refer to UAT2LLMs):
$MHA^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (6)
$FFN^{\prime}(\mathbf{x}^{\prime}_{i})=\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (7)
According to Eq. (6) and Eq. (7), the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq. (8) (see Appendix A for more details). The mathematical form of Eq. (8) can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq. (1). Therefore, if we parallelize multiple instances of $\mathbf{x}^{\prime}_{1}$, we obtain a multi-layer parallel Transformer network, as illustrated in Figure. 3.
Additionally, we provide the output for a two-layer Transformer, as shown in Eq. (9) (see Appendix A for more details). From Eq. (9), it is evident that the serial configuration of two Transformer layers also adheres to the mathematical form of UAT. Therefore, we can extend the Para-Former structure. Specifically, a block in Figure. 1 can contain multiple serial Transformer modules (it has been proven in UAT2LLMs that the mathematical form of multi-layer Transformers is also consistent with UAT), while multiple blocks operate in parallel, as illustrated in Figure. 4. The number of Transformer modules within a single block is referred to as the depth, while the total number of Transformer modules across all blocks is referred to as the network’s layer count.
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (9)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
Both design strategies possess the capability to dynamically adjust model parameters based on the input. In Figure 3, the primary mechanism is the dynamic adjustment of MHA weight parameters based on the input. In contrast, Figure 4 not only leverages the dynamic adjustment of MHA weight parameters but also utilizes UAT to dynamically approximate some bias parameters. For instance, in Eq. (9), $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$, can be seen as a two-layer UAT. Refer to Appendix B to see which parameters are influenced by the input. So, what are advantages of these strategies? We will talk about these in Section 4.3.

### v5

# Dynamic Universal Approximation Theory Foundation for Parallelism in CV Neural Networks

@@ Abstract
Neural networks are increasingly evolving towards training large models with big data, a method that has demonstrated superior performance across many tasks. However, this approach introduces an urgent problem: current deep learning models are serial computing, meaning that as the number of network layers increases, so do the training and inference times. This is unacceptable if deep learning models go deeper. Therefore, this paper proposes a deep learning parallelization strategy based on the Dynamic Universal Approximation Theorem (DUAT). From this foundation, we designed a parallel network called Para-Former to test our theory. Unlike traditional serial models, the inference time of Para-Former does not increase with the number of layers, significantly accelerating the inference speed of multi-layer networks. Experimental results validate the effectiveness of this network.
@@ 1 Introduction
Currently, deep learning is advancing toward larger-scale models and datasets, a trend driven by the successes of models such as ChatGPT Radford & Narasimhan (2018); Brown et al. (2020); Achiam et al. (2023), Llama Touvron et al. (2023), and ViT Dosovitskiy et al. (2020). These large models exhibit significantly enhanced performance compared to their predecessors. According to DUAT2LLMs Wang & Li (2024b) and DUAT2CVs Wang & Li (2024a), both residual-based convolutional neural networks (CNNs) and large-scale Transformer-based models are essentially DUAT functions Wang & Li (2024b; a). DUAT, an evolution of the Universal Approximation Theory (UAT) Cybenko (1989); Hornik et al. (1989), enables dynamic function approximation based on input data. From the DUAT perspective, larger models inherently possess a greater dynamic fitting capacity Wang & Li (2024b; a). Consequently, the shift toward larger models is undoubtedly the future of deep learning, and large datasets are essential for training these models to ensure their approximations accurately reflect real-world complexity.
@@
Current solutions include model optimization techniques such as quantization Jacob et al. (2017); Lu et al. (2022); Guo et al. (2021); Lin et al. (2023), pruning Ma et al. (2023); Xia et al. (2023), and knowledge distillation Gou et al. (2020); Gu et al. (2021). Additionally, specialized accelerators Jouppi et al. (2017); Elbtity et al. (2024) and edge computing Shi et al. (2016); Gale et al. (2020), as well as parallel and distributed computing Shallue et al. (2018); Ben-Nun & Hoefler (2018); Sergeev & Balso (2018), offer potential solutions. However, these methods do not fundamentally address the inference delay caused by the increasing number of serial network layers.
@@
The root cause lies in the serial computation mechanism of existing deep learning models, where the output of one layer serves as the input for the next. This design originated from early computer vision networks, which used convolutional layers with small receptive fields to accommodate the local correlations and spatiotemporal invariance of image data. Achieving a larger receptive field required deeper network layers, as exemplified by architectures like AlexNet Krizhevsky et al. (2012) and VGG Simonyan & Zisserman (2014). The introduction of ResNet He et al. (2015) marked a paradigm shift, as the residual structure could dynamically fit functions based on input Wang & Li (2024b; a), influencing subsequent network designs and establishing the tradition of multi-layer stacking. While this design was manageable with fewer layers, the increasing depth and number of parameters in modern networks have created a critical issue: slow computation efficiency due to the need for sequential layer-by-layer calculations during inference. This computation method inevitably leads to delays, which become more severe as network depth increases, significantly constraining the progress of computer vision and natural language processing technologies.
@@
Therefore, exploring parallelization techniques for deep learning networks—enabling parallel computation across multiple layers—is crucial for accelerating inference, overcoming time bottlenecks, and advancing computer vision and natural language processing technologies towards greater efficiency and speed. To develop such a parallel network, we build on UAT, integrating DUAT Wang & Li (2024b; a) to theoretically propose a foundational approach to parallel network design. To validate our theory, we designed a parallel network, Para-Former, based on the existing ViT architecture. Our contributions include:
  • We propose a parallel network approach based on DUAT. First, we outline the direction for parallel network design based on UAT and then identify the issues associated with this design. Next, we present the fundamental principles and requirements for parallel network design based on DUAT.
@@
  • We analyze the characteristics, advantages, and disadvantages of both serial and parallel networks from the perspective of DUAT.
@@
  • We investigate the issues on performance present in deep learning from a data perspective and give the solution.
The structure of this paper is as follows: In Section 2, we begin with a brief introduction to the general form of UAT (Section 2.1). Next, we present a method for designing parallel networks based on the UAT and discuss the prolems associated with such network designs (Section 2.2). To address these challenges, we introduce DUAT in Section 3, providing an initial overview in Section 3.1 before proposing the Para-former model based on DUAT in Section 3.2. Finally, we analyze the efficiency improvements of this parallel network in Section 3.3. In Section 4, we validate the effectiveness of Para-Former through various evaluations: assessing its feasibility in Section 4.2, examining the impact of network depth in Section 4.3, and analyzing the influence of the dataset on network performance in Section 4.4.
2 UAT to Parallel Network
2.1 Introduction of UAT
To develop a parallel network, we start with mathematical forms of UAT. Before formally designing the parallelized network, we will first review the generalized mathematical forms of UAT. The UAT was originally introduced by Cybenko, with the following basic form:
@@ 2.1 Introduction of UAT
This indicates that when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik et al. further proved that this function can approximate any Borel measurable function. Observing Eq. equation 1, where the function $G(\mathbf{x})$ produces scalar outputs in the real number field $\mathbb{R}$, the scenario naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation in each dimension. To accommodate such multi-dimensional outputs, we simply need to extend Eq. equation 1: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. We call the parameters $\mathbf{W}_{j}$ and $\alpha_{j}$ in UAT as weight and $\theta_{j}$ as bias.
2.2 The Problems in Designing parallel Network Based on UAT
Figure 1: The general description of the serial network and parallel network based on UAT.
@@ 2.2 The Problems in Designing parallel Network Based on UAT
From Eq. equation 1, we can observe a significant characteristic: if we define different indices $j$ as the indices of layers and the corresponding layer’s mathematical form is $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$, it is evident that the inputs, outputs and parameters between different layers do not interact with each other. Therefore, based on Eq. equation 1, we can design networks where parameters across multiple layers are independent, achieving parallel computation, like Figure 1. Figure 1.a represents a typical sequential network, which clearly requires computation to proceed in order from front to back. In contrast, Figure 1.b depicts a parallel network, where there is no parameter interaction between different blocks, allowing them to be considered independent and thus suitable for parallel computation. The blocks in this network should satisfy the formula of $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$. Clearly, under this condition, the sum of the multiple blocks’ outputs adheres to the UAT theory.
If we design the parallel network solely based on Eq. equation 1, the network will face a limitation: Bad performance. This network format is a single layer perception network MinskySeymour (1969) and it is out of date because of poor performance. The fundamental reason is that once training is complete: the function that can be approximated by the network is fixed, as the parameters in the corresponding UAT are fixed. Although UAT has powerful approximation ability, the real world is complex. It is impossible to approximate the world using a fix function.
3 DUAT to Parallel Network
In Section 2.2, we have given the drawback of designing parallel networks based on UAT straightforwardly. So to ensure the network can dynamically approximate the world, we adopt the DUAT. DUAT’s general mathematical form is the same as UAT, but some or all parameters in DUAT Wang & Li (2024b; a) can dynamically change based on input. This allows the network to dynamically adapt and fit the function based on the input.
3.1 Introduction of DUAT
The DUAT was introduced in Wang & Li (2024b; a), which can be seen as an advanced form of UAT based on the residual structure. To introduce DUAT, we first explain the Matrix-Vector Method Wang & Li (2024b; a). In neural networks, there are a variety of fundamental operations, such as convolution, pooling, multi-head attention (MHA), and feed-forward networks (FFN), they are represented primarily by engineer format rather than expressed in a unified mathematical form.
@@ 3.1 Introduction of DUAT
To address this, Wang & Li (2024b; a) introduced the Matrix-Vector Method, a technique that standardizes these operations within a unified mathematical framework, as shown in Figure 2. In this context, $\mathbf{x}$ and $\mathbf{y}$ denote the input and output of a basic operation, with $T(\mathbf{W})$ representing various transformations, such as convolution, and $\mathbf{W}$ corresponding to the parameters of the transformation $T$. The transformations $T_{D}$ and $T_{P}$ are applied to convert $\mathbf{x}$, $\mathbf{y}$, and $\mathbf{W}$ into column vectors $\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, and a parameter matrix $\mathbf{W}^{\prime}$, ensuring that the equation $\mathbf{W}^{\prime}\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$ is satisfied.
To clearly distinguish these transformed variables, we mark them with a prime symbol, for instance, $\mathbf{x}\mapsto\mathbf{x}^{\prime}$. With this standardization, we can express the operations of multilayer networks as a cohesive mathematical form, allowing for a unified representation across different network structures.
Figure 3: The process of Transformer module.
Following this approach, Figure 3 provides an example of the typical operations within a module in a Transformer model. According to the Matrix-Vector Method, MHA can be represented as Wang & Li (2024b):
$MHA(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}$ (3)
FFN can be expressed as Wang & Li (2024b):
$FFN(\mathbf{x}_{i})\mapsto\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}\$ (4)
According to Eq. equation 3 and Eq. equation 4, the output $\mathbf{x}^{\prime}_{1}$ of a single-layer Transformer can be expressed as Eq. equation 5 (see Appendix A for more details). For convenience, the mathematical form of Eq. equation 5 can be considered equivalent to the two terms $\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ in Eq. equation 1.
Using Figure 3 as a reference for a typical module in a multi-layer Transformer network, we can derive that a multi-layer Transformer is effectively a DUAT function, which can be expressed as:
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1}^{\prime})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (6)
For the case where $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$. For cases where $i\geq 1$, we define the parameter updates as follows. For each $j=1,2,\ldots,i$, the updates are: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,3}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}^{\prime}_{j,3}$. Additionally, for $j=2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (7)
$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
Besides, parameters within the MHA mechanism also vary dynamically with the input. Therefore, in the DUAT associated with Transformers, the parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input.
3.2 Para-former
In Section 3.1, we introduced DUAT using the Transformer model as an example. Following this rationale, we designed both single-depth and multi-depth parallel Transformer structures, as shown in Figures 4 and 5. We selected the Transformer for this design due to its robust performance and established utility across both CV and NLP domains.
@@ 3.2 Para-former
It is evident that Transformer-based models, whether single-layer or multi-layer, possess the capability to dynamically approximate functions based on input. According to Eq. equation 6, the parameters responsible for this dynamic approximation in a single layer are $\mathbf{W}^{\prime}_{1,1}$ and $\mathbf{W}^{\prime}_{1,2}$. In multi-layer structures, both weights and bias terms exhibit dynamic approximation capabilities. Notably, the bias terms also leverage DUAT for approximation. Consequently, the input-output relationship and parameters in a multi-layer Transformer structure are inherently interrelated.
The advantage of the single-depth model lies in its reduced computation time; however, only the weight terms are dynamically changed. In contrast, the multi-depth model benefits from dynamic approximation in both its weights and multiple bias terms, with bias terms also employing DUAT Wang & Li (2024b). These approaches in networks design raises two key questions: is this network design feasible, and if so, which network performs better when the number of basic modules is identical? We will address these questions in Section 4.3.
3.3 Speed-up Ratio
We have proposed a network parallelization scheme, and in this section, we will compare the running speed of this parallelized network with that of traditional serial networks. It is important to note that due to differences in hardware and implementation methods, providing an objective comparison through experiments can be challenging. Therefore, we will focus on a theoretical analysis. Based on the network structure shown in Figure 4 and 5, the inference time of Para-Former is only affected by its network depth.
4 Experiments
Our experimental approach is to first verify the feasibility of the parallel network theory and its adherence to the DUAT properties (The larger value of $N$, the more powerful the model becomes, resulting in better classification performance.) using several commonly used public image classification datasets (Section 4.2). We then explore the impact of network depth on the Para-Former performance (Section 4.3), followed by an investigation into the influence of the dataset (Section 4.4). We conducted experiments with various Para-Former models, where the naming convention uses the first number to indicate the depth and the second number to indicate the number of layers. For instance, Para-former-1-12 denotes a model with 1 depth and 12 layers.
@@ 4 Experiments
4.1 Dataset
In pursuit of a comprehensive understanding of the efficacy and practicality of parallel models, we meticulously selected several widely-acknowledged and representative image classification datasets for an integrated comparative analysis. These datasets encompass a diverse array of visual content and span various levels of difficulty, thereby furnishing us with a fertile ground to probe the performance boundaries of parallel models. Specifically, our study centers on the following pivotal datasets: CIFAR-10, CIFAR-100 Krizhevsky (2009), STL-10 Coates et al. (2011), Flowers-102 Nilsback & Zisserman (2008) and Oxford-IIIT Pets Parkhi et al. (2012). In Table 1, we provide more details for these datasets.